find the three geometric means between 3 and 3/16.?
Answer by stanbon(75887)

find the three geometric means between 3 and 3/16.?
---------
a(1) = 3
a(5) = a(1)*r^4 = 3/16
---
Solve:
r^4 = (3/16)/(3/1)
r^4 = 1/16
r = 1/2
----------------
Series:
3,3/2,3/4,3/8,3/16
---------
Cheers,
Stan H.
----------